taaron5241:

Line JK passes through points J(–3, 11) and K(1, –3). What is the equation of line JK in standard form? 7x + 2y = –1 7x + 2y = 1 14x + 4y = –1 14x + 4y = 1

surjithayer:

\[slope~m=\frac{ y_{2}-y_{1} }{ x_{2}-x_{1} }\] take either point J or K and then use the formula \[y-y_{1}=m(x-x_{1})\] and simplify